OpenStudy (anonymous):

{y=x-2 {y=4x+1 (These braces are just huge bit two)

OpenStudy (anonymous):

Ok so if y= x-2 and 4x+1 then it is safe to assume that x-2=4x+1 from there you need to get the Xs on one side and the numbers on the other to move a number to the other side you need to do the opposite of it. I decided bring the 1 over first x-2-1=4x+1-1 x-3=4x then the x x-x-3=4x-x -3=3x then divide -1=x